Use a system of equations to solve each problem
13. Adult tickets for the school musical sold for $3.50 and student tickets sold for $2.50. 321 tickets were sold altogether for $937.50. How many of each kind of ticket were sold?

Answer by ankor@dixie-net.com(22740)

Adult tickets for the school musical sold for $3.50 and student tickets sold for $2.50. 321 tickets were sold altogether for $937.50. How many of each kind of ticket were sold?
:
Let a = no. of adult tickets
Let s = no. of student tickets
:
It said "a total of 321 tickets were sold", therefore:
a + s = 321
or
s = (321 - a)
:
It said total revenue was 937.50
3.5a + 2.5s = 937.50
:
In the above equation, substitute (321-a) for s; find a:
3.5a + 2.5(321-a) = 937.50
:
3.5a + 802.5 - 2.5a = 937.50
:
3.5a - 2.5a = 937.5 - 802.5
:
a = 135 adults
:
s = 321 - 135 = 186 students
:
:
Check our solutions:
3.5(135) + 2.5(186) =
472.5 + 465 = 937.50